# Deep Active Learning in the Open World

### v1

@@ 5.2.2 Main Results on Multiple Datasets
We now present the results of our algorithm applied to CIFAR100-LT, ImageNet-LT, and Places365-LT. As mentioned before, the long-tailed distribution of classes is a well-known challenge in practical applications. Due to this issue, we often encounter open world learning scenarios where rare classes are likely to be missed during the initial annotation phase. In the following experiments, we receive annotations from only a small number of classes in the initial batch of annotation (three for CIFAR100-LT and Places365-LT, and ten for ImageNet-LT). In addition, our algorithm ALOE use energy and GradNorm for OOD scoreing function $\Omega$ and k-means for clustering method. We also include ablation studies around different initial number of classes, OOD scores, and clustering methods in the later sections.
(a) CIFAR100-LT
(b) ImageNet-LT
(c) Places365-LT
Figure 3: Balanced test accuracy and number of annotated classes on CIFAR100-LT, ImageNet-LT and Places365-LT. Reverse ALOE result shown on CIFAR100-LT. (Shaded region represent the standard error conducted across four trials.)
@@
As shown in Figure 3, on CIFAR100-LT, our method ALOE significantly outperforms all baseline methods in both balanced accuracy and the number of newly discovered classes. Starting with only three initially labeled classes, our method quickly expands the set of annotated classes, maintaining a higher discovery rate throughout the iterations. By leveraging the combination of OOD detection and diversity-based sampling, ALOE is able to efficiently explore unknown classes while ensuring strong performance on the known classes. On ImageNet-LT, ALOE saves 70% of annotation cost to achieve the same accuracy comparing to random sampling. We also note that while Coreset and Galaxy occasionally match ALOE’s performance, they struggle with consistency across different datasets. In contrast, ALOE consistently performs best among all methods on all datasets.
@@
We also experimented with Reverse ALOE, which doesn’t show sufficient label efficiency comparing to other baselines. The Reverse ALOE result shown in Figure 3 uses the combination of energy OOD score and Coreset clustering method, which behaves best in all combinations. While still competitive with some baselines, Reverse ALOE does not match the performance of ALOE. This performance gap likely stems from conducting OOD filtering first, which eliminates many classes of out-of-distribution examples early in the process. As a result, Reverse ALOE works with a less diverse set of examples compared to ALOE, which preserves diversity by clustering examples before performing OOD filtering.
@@ 5.2.4 Ablation Study of OOD Scores
Figure 5 shows that different OOD scores affect the performance of our algorithm. On CIFAR100-LT, Energy, MSP and GradNorm scores perform similarly. While MSP and Energy scores both focus on predictions, Energy’s consideration of all class predictions makes it potentially more resilient on complex datasets. Thus, Energy and GradNorm scores are used in main experiments 5.2.2. Energy score is used for CIFAR100-LT and ImageNet-LT, while GradNorm is applied to Places365-LT.
Figure 5: Ablation experiment for using different OOD scores $\Omega(x,f)$ in ALOE on CIFAR100-LT. (Shaded region represent the standard error conducted across four trials.)
@@ 5.2.5 Ablation Study of Clustering Methods
We also investigated various clustering methods. We experimented with popular clustering method used in prior active learning algorithms. For instance, TypiClust (Hacohen et al., 2022) employs k-means, Coreset (Sener & Savarese, 2017) utilizes k-center, and BADGE (Ash et al., 2019) uses k-means++. As shown in Figure 6, k-means yielded the best results, and we adopted it for all other experiments.
Figure 6: Ablation experiment for clustering methods on CIFAR100-LT with energy OOD score. (Shaded region represent the standard error conducted across four trials.)
@@ 6 Balancing Novel Class Discovery and Known Class Learning
When running ALOE further on a larger amount of annotation budget, our experiments (Figure 7) demonstrate an intriguing phenomenon – baseline methods can even achieve slightly better performance than ALOE once most classes have been identified. Similarly, as mentioned in 5.2.3, on CIFAR100-LT, when the number of initial classes increases, the improvement gap of using ALOE narrows. These findings are expected, as ALOE uniquely prioritizes annotating diverse OOD classes, while baseline approaches focus on learning known ID classes. While this suggests the potential for an algorithm that excels at both tasks, achieving this dual objective presents significant challenges.

### v2

@@ 5.2.2 Main Results on Multiple Datasets
We now present the results of our algorithm applied to CIFAR100-LT, ImageNet-LT, and Places365-LT. As mentioned before, the long-tailed distribution of classes is a well-known challenge in practical applications. Due to this issue, we often encounter open world learning scenarios where rare classes are likely to be missed during the initial annotation phase. In the following experiments, we receive annotations from only a small number of classes in the initial batch of annotation (three for CIFAR100-LT and Places365-LT, and ten for ImageNet-LT). In addition, our algorithm ALOE uses GradNorm for OOD scoreing function $\Omega$ and k-means for clustering method. We also include ablation studies around different initial number of classes, OOD scores, and clustering methods in the later sections. In Table 3 and Table 4, balanced test accuracy on the three datasets are shown. Number of annotated classe and the plot of the results are summarized in Table 5, Table 6 and Figure  8 in Appendix B.
Table 3: Balanced test accuracy on CIFAR100-LT and Places365-LT with different budget. (The standard error was calculated based on four trails.)
Dataset CIFAR100-LT Places365-LT
Budget $250$ $750$ $600$ $2000$
Random $0.392\small{\scriptstyle\pm 0.006}$ $0.553\scriptstyle\pm 0.004$ $0.208\scriptstyle\pm 0.007$ $0.275\scriptstyle\pm 0.004$
Margin $0.445\scriptstyle\pm 0.006$ $0.643\scriptstyle\pm 0.011$ $0.217\scriptstyle\pm 0.005$ $0.287\scriptstyle\pm 0.004$
Badge $0.447\scriptstyle\pm 0.006$ $\mathbf{0.652\scriptstyle\pm 0.008}$ $0.231\scriptstyle\pm 0.004$ $0.295\scriptstyle\pm 0.004$
Galaxy $0.436\scriptstyle\pm 0.007$ $0.645\scriptstyle\pm 0.008$ $0.231\scriptstyle\pm 0.004$ $\mathbf{0.302\scriptstyle\pm 0.005}$
Coreset $0.449\scriptstyle\pm 0.012$ $0.617\scriptstyle\pm 0.010$ $0.238\scriptstyle\pm 0.005$ $\mathbf{0.304\scriptstyle\pm 0.003}$
ProbCover $0.346\scriptstyle\pm 0.010$ $0.442\scriptstyle\pm 0.004$ $0.152\scriptstyle\pm 0.007$ $0.247\scriptstyle\pm 0.005$
TypiClust $0.376\scriptstyle\pm 0.015$ $0.509\scriptstyle\pm 0.004$ $0.199\scriptstyle\pm 0.005$ $0.275\scriptstyle\pm 0.004$
EOAL $0.368\scriptstyle\pm 0.005$ $0.578\scriptstyle\pm 0.005$ $0.211\scriptstyle\pm 0.007$ $0.277\scriptstyle\pm 0.005$
ALOE(Ours) $\mathbf{0.479\scriptstyle\pm 0.007}$ $\mathbf{0.650\scriptstyle\pm 0.04}$ $\mathbf{0.254\scriptstyle\pm 0.004}$ $\mathbf{0.310\scriptstyle\pm 0.008}$
Table 4: Balanced test accuracy on ImageNet-LT with different budget. (The standard error was calculated based on four trails.)
Dataset ImageNet-LT
Budget $3000$ $10000$
Random $0.391\scriptstyle\pm 0.002$ $0.541\scriptstyle\pm 0.003$
Margin $0.424\scriptstyle\pm 0.006$ $0.562\scriptstyle\pm 0.002$
Badge $0.434\scriptstyle\pm 0.004$ $0.581\scriptstyle\pm 0.002$
Galaxy $0.438\scriptstyle\pm 0.006$ $\mathbf{0.631\scriptstyle\pm 0.001}$
Coreset $0.351\scriptstyle\pm 0.005$ $0.525\scriptstyle\pm 0.002$
ALOE(Ours) $\mathbf{0.517\scriptstyle\pm 0.004}$ $\mathbf{0.617\scriptstyle\pm 0.003}$
@@
As shown in Table 3 and Table 4, our method ALOE significantly outperforms all baseline methods in both balanced accuracy and the number of newly discovered classes. Starting with only three initially labeled classes, our method quickly expands the set of annotated classes, maintaining a higher discovery rate throughout the iterations. By leveraging the combination of OOD detection and diversity-based sampling, ALOE is able to efficiently explore unknown classes while ensuring strong performance on the known classes. On ImageNet-LT, ALOE saves 70% of annotation cost to achieve the same accuracy comparing to random sampling. We also note that while some baselines occasionally match ALOE’s performance, they struggle with consistency across different datasets. In contrast, ALOE consistently performs best among all methods on all datasets.
@@
We also experimented with Reverse ALOE, which doesn’t show sufficient label efficiency comparing to other baselines. The Reverse ALOE result shown in Figure 8 uses the combination of GradNorm OOD score and Coreset clustering method, which behaves best in all combinations. While still competitive with some baselines, Reverse ALOE does not match the performance of ALOE. This performance gap likely stems from conducting OOD filtering first, which eliminates many classes of out-of-distribution examples early in the process. As a result, Reverse ALOE works with a less diverse set of examples compared to ALOE, which preserves diversity by clustering examples before performing OOD filtering.
@@ 5.2.4 Ablation Study of OOD Scores
Figure 4 shows that different OOD scores affect the performance of our algorithm. On CIFAR100-LT, Energy and GradNorm scores perform similarly. However, we used GradNorm to carry out our main experiments in Section 5.2.2, since it is a direct reflection on the uncertainty of the model, espiecially deep learning models.
Figure 4: Ablation experiment for using different OOD scores $\Omega(x,f)$ in ALOE on CIFAR100-LT. (Shaded region represents the standard error conducted across four trials.)
@@ 5.2.5 Ablation Study of Clustering Methods
We also investigated various clustering methods. We experimented with a popular clustering method used in prior active learning algorithms. For instance, TypiClust (Hacohen et al., 2022) employs k-means, Coreset (Sener & Savarese, 2017) utilizes k-center, and BADGE (Ash et al., 2019) uses k-means++. We also implemented gaussian mixture models and mini-batch K-means as the clustering subprocedure. Both of these are more computationally efficient than simple k-means, and obtain similar performance as the procedure using k-means. As shown in Figure 5, this indicates that our algorithm is not highly sensitive to the choice of clustering algorithm. In our implementation, we used k-means despite its relatively higher computational cost compared to other clustering methods. However, in most active learning scenarios, the selection cost is significantly lower than the cost of training the neural network. In the rare cases where k-means becomes impractical, we recommend practitioners consider alternatives such as Gaussian Mixture Models (GMM) or mini-batch k-means.
Figure 5: Ablation experiment for clustering methods on CIFAR100-LT with GradNorm OOD score. (Shaded region represent the standard error conducted across four trials.)
@@ 6 Balancing Novel Class Discovery and Known Class Learning
When running ALOE further on a larger amount of annotation budget, our experiments (Figure 6) demonstrate an intriguing phenomenon – baseline methods can even achieve slightly better performance than ALOE once most classes have been identified. Similarly, as mentioned in Section 5.2.3, on CIFAR100-LT, when the number of initial classes increases, the improvement gap of using ALOE narrows. These findings are expected, as ALOE uniquely prioritizes annotating diverse OOD classes, while baseline approaches focus on learning known ID classes. While this suggests the potential for an algorithm that excels at both tasks, achieving this dual objective presents significant challenges.
@@ 7 Conclusion
Acknowledgement
This work has been supported by ONR MURI grant N00014-20-1-2787.
@@ Appendix A Long Tail Datasets
Appendix B Experiments Results
(a) CIFAR100-LT
(b) ImageNet-LT
(c) Places365-LT
Figure 8: Balanced test accuracy and number of annotated classes on CIFAR100-LT, ImageNet-LT and Places365-LT with all baselines and our algorithms. Reverse ALOE result shown on CIFAR100-LT. (Shaded region represent the standard error conducted across four trials.)
Table 5: Number of annotated classes on CIFAR100-LT and Places365-LT with different budget. (The standard error was calculated based on four trails.)
Dataset CIFAR100-LT Places365-LT
Budget $250$ $750$ $600$ $2000$
Random $57.75\scriptstyle\pm 0.63$ $81.75\scriptstyle\pm 1.11$ $188.00\scriptstyle\pm 6.36$ $275.50\scriptstyle\pm 3.30$
Margin $66.75\scriptstyle\pm 1.31$ $94.75\scriptstyle\pm 1.31$ $203.75\scriptstyle\pm 4.61$ $287.75\scriptstyle\pm 2.02$
Badge $66.75\scriptstyle\pm 0.75$ $\mathbf{96.00\scriptstyle\pm 0.82}$ $209.75\scriptstyle\pm 2.87$ $295.25\scriptstyle\pm 3.38$
Galaxy $68.50\scriptstyle\pm 2.50$ $94.50\scriptstyle\pm 0.29$ $207.25\scriptstyle\pm 7.08$ $\mathbf{306.50\scriptstyle\pm 2.33}$
Coreset $66.25\scriptstyle\pm 1.70$ $91.0\scriptstyle\pm 0.71$ $\mathbf{221.25\scriptstyle\pm 2.53}$ $\mathbf{312.75\scriptstyle\pm 1.93}$
ProbCover $51.50\scriptstyle\pm 0.96$ $61.75\scriptstyle\pm 0.63$ $132.25\scriptstyle\pm 8.15$ $228.00\scriptstyle\pm 5.05$
TypiClust $56.00\scriptstyle\pm 1.08$ $73.25\scriptstyle\pm 1.03$ $181.00\scriptstyle\pm 4.64$ $265.00\scriptstyle\pm 2.48$
EOAL $53.75\scriptstyle\pm 0.95$ $85.5\scriptstyle\pm 1.32$ $182.25\scriptstyle\pm 1.80$ $256.00\scriptstyle\pm 3.92$
ALOE(Ours) $\mathbf{73.5\scriptstyle\pm 0.94}$ $\mathbf{96.1\scriptstyle\pm 0.51}$ $\mathbf{230.50\scriptstyle\pm 1.80}$ $\mathbf{304.25\scriptstyle\pm 2.17}$
Table 6: Number of annotated classes on ImageNet-LT with different budget. (The standard error was calculated based on four trails.)
Dataset ImageNet-LT
Budget $3000$ $10000$
Random $584.75\scriptstyle\pm 2.90$ $847.75\scriptstyle\pm 5.02$
Margin $639.50\scriptstyle\pm 10.73$ $871.00\scriptstyle\pm 1.96$
Badge $650.75\scriptstyle\pm 4.50$ $915.25\scriptstyle\pm 3.12$
Galaxy $660.50\scriptstyle\pm 12.84$ $\mathbf{989.0\scriptstyle\pm 1.29}$
Coreset $536.75\scriptstyle\pm 5.02$ $824.25\scriptstyle\pm 5.19$
ALOE(Ours) $\mathbf{789.25\scriptstyle\pm 4.66}$ $\mathbf{977.00\scriptstyle\pm 2.04}$
Appendix C Per-batch Running Time Analysis of ALOE and Baselines
We analyze the per-batch running time of ALOE in comparison to the mentioned baselines. Recall that $B$ is the batch size, $N$ is the pool size, $K$ is the number of classes, $Q$ is the forward inference time of a neural network on a single example, and $d$ is the feature dimension. Random Sampling is computationally the simplest, with a time complexity of $O(B)$, as examples are chosen uniformly at random without any additional computation. Confidence Sampling methods, such as Margin Sampling and EOAL, require $O(QN+KN+B\log N)$, where $O(QN)$ comes from forward passes on the pool, $O(KN)$ from computing uncertainty scores (e.g., margins or entropy), and $O(B\log N)$ from selecting the top $B$ uncertain examples.
Diversity-focused methods like Badge and Coreset involve more computational overhead. Badge combines uncertainty and diversity through gradient embeddings, requiring $O(QN+Nd^{2}+Bd^{3})$, where $O(Nd^{2})$ arises from gradient embedding computation and $O(Bd^{3})$ from k-means++ clustering. Coreset solves a k-center problem, leading to a complexity of $O(QN+N^{2}d)$ in its exact form, although approximation methods reduce this to $O(QN+BNd)$. TypiClust is simpler, with $O(QN+BN)$, and ProbCover incurs $O(QN+N^{2}d)$ due to pairwise coverage probability computations.
For active learning algorithms in practice, time complexity is often not a critical bottleneck because $O(QN)$, representing the cost of forward passes on the pool, dominates all other terms. This ensures that methods like ALOE remain computationally efficient while incorporating both diversity and OOD detection to achieve superior performance in open-world active learning scenarios.